# FakET: Simulating Cryo-Electron Tomograms with Neural Style Transfer

### v1

@@ Abstract
Particle localization and -classification constitute two of the most fundamental problems in computational microscopy. In recent years, deep learning based approaches have been introduced for these tasks with great success. A key shortcoming of these supervised learning methods is their need for large training data sets, typically generated from particle models in conjunction with complex numerical forward models simulating the physics of transmission electron microscopes. Computer implementations of such forward models are computationally extremely demanding and limit the scope of their applicability. In this paper we propose a simple method for simulating the forward operator of an electron microscope based on additive noise and Neural Style Transfer techniques. We evaluate the method on localization and classification tasks using one of the established state-of-the-art architectures showing performance on par with the benchmark. In contrast to previous approaches, our method accelerates the data generation process by a factor of 750 while using 33 times less memory and scales well to typical transmission electron microscope detector sizes. It utilizes GPU acceleration and parallel processing. It can be used as a stand-alone method to adapt a training data set or as a data augmentation technique. The source code is available at https://gitlab.com/deepet/faket.
Keywords— Machine Learning, Deep Learning, CryoET, Cryo-Electron Tomography, Neural Style Transfer, CryoEM, Deep Finder, Synthetic Data, Data Augmentation, Pre-training, Domain Adaptation, Classification, Localization
Acknowledgement— The IMP and D.H. are generously funded by Boehringer Ingelheim. We thank Julius Berner from the Mathematical Data Science @ UniVie, Ilja Gubins and Marten Chaillet from the SHREC team, and the members of the Haselbach lab for helpful discussions.
Correspondence to— P. Harar pavol.harar@univie.ac.at
@@ 1 Introduction
| Model | Training data | Data cost | Localization F1 | Classification F1 macro |
@@
Recent developments in cryo-electron tomography (cryoET) allow to obtain high resolution representations of macromolecular complexes in their native cellular environment showing molecular interactions that are hardly accessible with other methods Turk and Baumeister [2020]. In cryoET, the imaged sample is in most cases a 100-200 nm thick slice of a frozen cell. From this slice, projection images are taken in a transmission electron microscope (TEM) from different rotation (tilt) angles. An artifact free reconstruction would require measurements using tilt angles, that would complete the half circle. However, this is not feasible, due to limitations of the specimen holder, and only a range of 140∘ can be recorded. The missing tilt images later on result in a so-called missing-wedge in the 3D reconstruction (cryo-electron tomogram). In addition, the electron beam severely damages the sample during imaging, so only a low electron dose can be used to image a biological specimen. The low dose in combination with the presence of ice in the sample results in the acquired data being very noisy. Consequently, the identification of molecules within these reconstructions is a daunting task. Particle identification is however necessary as the particles need to be classified and averaged to determine high resolution structures. While cryoET has led to a large number of breakthroughs, providing hitherto unseen detail in the molecular architecture of cells Zimmerli et al. [2021], O’Reilly et al. [2020], Mahamid et al. [2016], the aforementioned challenges still hinder the widespread use of cryoET in the larger cell biology and structural biology community. In this context, the development of new reliable software tools is of paramount importance, which is however obstructed by the lack of sufficient accessible and annotated data to develop the software tools on.
@@ 1.1 SHREC simulator
To overcome the problem with the lack of data, in 2019, the annual SHREC – 3D Shape Retrieval Contest included a new track titled Classification in Cryo-Electron Tomograms. The organizers of this track proposed a task of localization and classification of biological particles in cryo-electron tomograms. Every year since, experts from 3D object retrieval and 3D electron microscopy communities were invited to participate in the challenge. In order to ensure fair evaluation and comparable results across the submissions, the organizers created a data set of ten physics-based cryo-electron tomogram simulations for the contestants to train and evaluate their methods on. Each year, the results of the contesting methods were presented and compared Gubins et al. [2019, 2020, 2021].
@@ 1.2 Our contribution
In this paper, we propose faket, a novel method for simulating the forward operator of a cryo transmission electron microscope to generate synthetic tilt-series. It was created, among other reasons, to generate training data for deep neural networks to solve tasks such as particle localization and (much more challenging) particle classification. Our method combines additive noise and neural style transfer (NST) techniques based on Gatys et al. [2016] and is evaluated on downstream tasks using DeepFinder (DF) Moebel et al. [2021] – a neural network specifically tailored to the task of particle localization and classification, see also Section 4.1. faket is a surrogate model that mimics the behavior of the physics-based TEM simulator SHREC, while being drastically less computationally expensive, both in terms of time and memory (see Table 1). It also does not require any calibration protocol as other simulators, see Section 1.4. Moreover, it does not need to be retrained to be used on new data, nor does it require labeled reference data. In contrast to SHREC, our method accelerates the data generation process by a factor of 750 while using 33 times less memory. It therefore has the potential to save experts countless hours 111Depending on the imaged sample, an expert may spend several hours of manual work per tilt-series per particle labeling the data. At the same time, certain smaller particles cannot be visually found at all, thus fully labeled tomograms do not really exist. of manual work in labeling their data sets. Moreover, our method is capable of simulating large tilt-series, which are common in experimental environments (about larger than SHREC projections). For example, we generated a tilt-series on a single NVIDIA A100 40GB SXM4 GPU in less than 10 minutes. This advance makes it possible to train particle localization and classification networks from scratch, or to pre-train networks that are later fine-tuned using manually labelled experimental data. On top of that, our method is open source and our experiments are reproducible. This paper also includes comparisons of the benchmark and the proposed faket method with a simple baseline method based on additive Gaussian noise, which will be useful for practitioners who need to decide which method fits into their computational budget.
@@ 1.3 Related work
A similar idea in X-ray-based computed tomography angiography (CTA) was investigated in Seemann et al. [2020]. The authors focused on solving a lumen segmentation task. However, due to the very different nature of the samples imaged in CTA 222The objects of interest in CTA are on average two orders of magnitude larger in relation to the size of the tomogram than those in cryoET, where the particles often span only tens of voxels., it was not clear whether a similar NST based framework could be successfully applied in cryoET to image nanoscale particles. In addition, the article does not provide source code and does not document the experiments in enough detail for us to be able to reproduce the results or adapt the method to the cryoET domain.
@@ 1.4 Other simulators
The value of simulated data in cryoEM is well recognized and micrograph simulation has been attempted several times. In first approximation a single cryoEM micrograph is the projection of a 3D object, convolved with the electron microscope’s point spread function. Additionally, the overall process comprises several sources of noise coming from the nature of the sample, the microscope, and the imaging procedure which are hard to model accurately. For a better overview of the attempts, it is necessary to mention TEM simulators that were developed in the past decades to simulate micrographs in cryo conditions. Earlier works provided fast simplistic models mostly based on additive white Gaussian noise333This motivated our choice of baseline. or coloured noise Scheible et al. [2021]. Other works originated from the insights into the physics of TEM image formation and advanced the simulations by modeling various sources of noise, e.g. an improvement in modeling the structural noise was proposed in TEM Simulator (C) Rullgård et al. [2011]. Another improvement was done in InSilicoTEM (Matlab) presented in Vulović et al. [2013] by implementing the multislice method originally proposed in Cowley and Moodie [1957]. This makes it the next most relevant simulator related to our work after SHREC which is, as many others, also based on the same multislice method 444This motivated our choice of benchmark.. To the best of our knowledge, the most recent improvement presented in Himes and Grigorieff [2021] is distributed as a part of cisTEM package (C++). It introduced frozen plasmon method to explicitly model spatially variable inelastic scattering processes in cryo-electron microscopy. The aforementioned works, based on the same multislice method, however, suffer from a heavy computational burden or are limited to simulations of a single molecular complex. Despite these limitations, they represent a set of precise advanced physics-based simulators and as such also provide a comprehensive literature survey referring the reader to a body of detailed resources about modeling the image formation in transmission electron microscopy. In material sciences, GPU accelerated simulators such as MULTEM (C++, CUDA) Lobato and Van Dyck [2015], Lobato et al. [2016], abTEM (Python) Madsen and Susi [2020], or Prismatic (C++, CUDA) Pryor et al. [2017] have emerged. However, as reviewed in Kirkland [2020], advanced TEM simulators used in material sciences require atomic models of the background and the entire specimen. Moreover, in cryoEM the sample and its interaction with the electron beam is fundamentally different. While in material science samples are often only a few layers of atoms thick and the imaged atoms have strong interaction with the electron beam, cryoEM samples are thousands of atom layers thick and are less ordered and only weakly interact with the electron beam. Additionally, cryoEM samples can only withstand little radiation before complete destruction resulting in much lower signal to noise ratios. Due to these differences, such simulators are of limited use for macromolecular biological specimen simulations that we experiment with in this article. Finally, in contrary to our approach, all aforementioned methods require precise calibration protocols for setting the values of simulation parameters.
@@ 1.5 Structure of this article
To evaluate the methods, we used the simulated data from each of them as a training set for the DeepFinder neural network (more in Section 4.1). The network was trained to solve two evaluation tasks proposed in SHREC challenge. The final comparison was done by observing the models’ performances on the SHREC test tomogram. Detailed description of the evaluation is in Section 4. More on experiments and results in Section 5. This paper is accompanied with source code, the results are fully-reproducible, and the full experiment results are available in the repository.
@@ 2 Data
For the purposes of this paper, we used the latest version of the SHREC data set from the year 2021. More specifically, we based the presented methods on full-resolution grandmodels (synthetic volumes containing randomly scattered particles) and simulated projections from the SHREC 2021 additional data set that was made publicly available555Available for download at: shrec.net/cryo-et. An example visualization of a simulated projection using the SHREC simulator is depicted in Figure 1. We use the unchanged simulated projections as a benchmark which we try to approximate with our proposed method.
@@
The steps to create the simulations were described in Gubins et al. [2021]. However, at the time of writing of this paper, the implementation is not publicly available and from the description of the steps it is clear that the method is, and rightfully so, very elaborate, and not at all elementary to reimplement (also due to various steps being under-documented, see Appendix B).
@@
On the following lines, we briefly summarize the steps needed to create the SHREC data set. The authors first constructed 3D ground-truth specimens (grandmodels) along with annotations. Each specimen contained uniformly distributed and rotated protein instances (represented as Coulomb density) of varying size and structure from the Protein Data Bank (PDB) Berman et al. [2000], as well as membranes and gold fiducials, which are commonly found in tomograms. Then, they simulated a layer of amorphous ice into the grandmodels before rotating them over 61 evenly-spaced tilt angles, ranging from -60∘ to +60∘, in order to be projected. The noiseless projections were produced using their own implementation of a TEM simulator based on the multislice approach presented in Vulović et al. [2013]. Next, they sampled from a Poisson distribution with a specific electron dose to obtain the final electron counts (files labeled as projections). After that, they randomly shifted the projections to model the tilt misalignment and scaled the amplitudes in Fourier space using information about amplitudes from experimental images in order to increase their similarity (files labeled as projections_unbinned). In the end, they obtained the final images (files labeled as reconstruction) by reconstructing the binned projections using weighted-backprojection algorithm from a private version of the PyTom package Hrabe et al. [2012].
@@
We chose the SHREC data set as it is established, well-executed, downloadable, and allows researchers to compare their results with previous works without the need to reproduce all results from scratch. In this paper, we are not focusing on solving the proposed tasks, instead, we are using them as evaluation metrics to measure the performance of our proposed methods. However, the availability of the data and descriptions of the methods made this research feasible.
@@
As stated earlier, we used the full-resolution grandmodels from the SHREC data set to create our own noiseless projections using Radon transform, as implemented in the scikit-image library for image processing in Python Van der Walt et al. [2014]. There are two reasons why we created our own noiseless projections. First being simply the fact that SHREC provided only the full-resolution noiseless projections already embedded in the simulated layer of ice, which was the behaviour we also wanted to include in our surrogate. Second being the fact that we wanted to make the full implementation of all our steps available in the code-base. We also had to omit the usage of the reconstructions provided by SHREC because the exact configuration of all the steps is not published and the version of the PyTom software package used to create this data is not public. Also, the public version is not yet properly documented or straightforward to use. Thus, the only feasible option to ensure comparable results across our experiments was to create our own reconstructions from all the newly created projections described in Section 3. Nevertheless, we put a lot of effort into matching our reconstructions with the ones originally published in SHREC. We created them using the RadonTea package Müller, Paul [2013] using a custom filter described in Appendix B. However, we used the original SHREC model_9 reconstruction as a test tomogram for all presented methods as it was done in all the challenges.

### v2

@@ Abstract
Particle localization and -classification constitute two of the most fundamental problems in computational microscopy. In recent years, deep learning based approaches have been introduced for these tasks with great success. A key shortcoming of these supervised learning methods is their need for large training data sets, typically generated from particle models in conjunction with complex numerical forward models simulating the physics of transmission electron microscopes. Computer implementations of such forward models are computationally extremely demanding and limit the scope of their applicability. In this paper we propose a method for simulating the forward operator of an electron microscope based on additive noise and Neural Style Transfer techniques. We evaluate the method on localization and classification tasks using one of the established state-of-the-art architectures showing performance on par with the benchmark. In contrast to previous approaches, our method accelerates the data generation process by a factor of 750 while using 33 times less memory and scales well to typical transmission electron microscope detector sizes. It utilizes GPU acceleration and parallel processing. It can be used to adapt a synthetic training data set according to reference data from any transmission electron microscope. The source code is available at https://gitlab.com/deepet/faket.
Keywords— Machine Learning, Deep Learning, CryoET, Cryo-Electron Tomography, Neural Style Transfer, CryoEM, Deep Finder, Synthetic Data Generation, Pre-training, Domain Adaptation, Classification, Localization
Acknowledgement— The IMP and D.H. are generously funded by Boehringer Ingelheim. We thank Julius Berner from the Mathematical Data Science group @ UniVie, Ilja Gubins and Marten Chaillet from the SHREC team, and the members of the Haselbach lab for helpful discussions.
Correspondence to— P. Harar pavol.harar@univie.ac.at and D. Haselbach david.haselbach@imp.ac.at
@@ 1 Introduction
| Model | Training data | Data cost | Localization F1 | Classification F1 |
@@
| DeepFinder | faket + fine-tuning | faket cost + manual labeling | 0.821 | 0.565 () |
@@
Recent developments in cryo-electron tomography (cryoET) allow to obtain high resolution representations of macromolecular complexes in their native cellular environment showing molecular interactions that are hardly accessible with other methods Turk and Baumeister (2020). In cryoET, the imaged sample is in most cases a 100-200 nm thick slice of a frozen cell. From this slice, projection images are taken in a transmission electron microscope (TEM) from different rotation (tilt) angles. An artifact free reconstruction would require measurements using tilt angles, that would complete the half circle. However, this is not feasible, due to limitations of the specimen holder, and only a range of 140∘ can be recorded. The missing tilt images later on result in a so-called missing-wedge in the 3D reconstruction (cryo-electron tomogram). In addition, the electron beam severely damages the sample during imaging, so only a low electron dose can be used to image a biological specimen. The low dose in combination with the presence of ice in the sample results in the acquired data being very noisy. Consequently, the identification of molecules within these reconstructions is a daunting task. Particle identification is however necessary as the particles need to be classified and averaged to determine high resolution structures. While cryoET has led to a large number of breakthroughs, providing hitherto unseen detail in the molecular architecture of cells (Zimmerli et al., 2021; O’Reilly et al., 2020; Mahamid et al., 2016), the aforementioned challenges still hinder the widespread use of cryoET in the larger cell biology and structural biology community. In this context, the development of new reliable software tools is of paramount importance, which is however obstructed by the lack of sufficient accessible and annotated data to develop the software tools on.
@@ 1.1 SHREC simulator
To overcome the problem with the lack of data, in 2019, the annual SHREC – 3D Shape Retrieval Contest included a new track titled Classification in Cryo-Electron Tomograms. The organizers of this track proposed a task of localization and classification of biological particles in cryo-electron tomograms. In the following years, experts from 3D object retrieval and 3D electron microscopy communities were invited to participate in the challenge. In order to ensure fair evaluation and comparable results across the submissions, the organizers created a data set of ten physics-based cryo-electron tomogram simulations (9 train & 1 test tomogram) for the contestants to train and evaluate their methods on. Each year, the results of the contesting methods were presented and compared Gubins et al. (2019, 2020, 2021).
@@ 1.2 Our contribution
In this paper, we propose faket, a novel method for simulating the forward operator of a cryo transmission electron microscope to generate synthetic tilt-series. It was created, among other reasons, to generate training data for deep neural networks to solve tasks such as particle localization and (much more challenging) particle classification. Our method combines additive noise and neural style transfer (NST) techniques based on Gatys et al. (2016) and is evaluated on downstream tasks using DeepFinder (DF) Moebel et al. (2021) – a neural network specifically tailored to the task of particle localization and classification, see also Section 4.1.
faket is a surrogate model that mimics the behavior of the physics-based TEM simulator SHREC, while being drastically less computationally expensive, both in terms of time and memory (see Table 1). It also does not require any calibration protocol as other simulators, see Section 1.4. Moreover, it does not need to be retrained to be used on new data, nor does it require labeled reference data. In contrast to SHREC, our method accelerates the data generation process by a factor of 750 while using 33 times less memory. It therefore has the potential to save experts countless hours 111Depending on the imaged sample, an expert may spend several hours of manual work per tilt-series per particle labeling the data. At the same time, certain smaller particles cannot be visually found at all, thus fully labeled tomograms do not really exist. of manual work in labeling their data sets.
Moreover, our method is capable of simulating large tilt-series, which are common in experimental environments (about larger than SHREC projections). For example, we generated a tilt-series on a single NVIDIA A100 40GB SXM4 GPU in less than 10 minutes. This advance makes it possible to train particle localization and classification networks from scratch, or to pre-train networks that are later fine-tuned using manually labelled experimental data, see Appendix C. On top of that, our method is open source and our experiments are reproducible.
Additionally, we include comparisons of the benchmark and the proposed faket method with a simple baseline method based on additive Gaussian noise, which will be useful for practitioners who need to decide which method fits into their computational budget. The contribution of our method is further supported by an ablation study presented in Appendix D, where we also offer insight into the potential limits of DF on the studied data by conducting experiments using completely noiseless simulated tomograms.
@@ 1.3 Related work
A similar idea in X-ray-based computed tomography angiography (CTA) was investigated in Seemann et al. (2020). The authors focused on solving a lumen segmentation task. However, due to the very different nature of the samples imaged in CTA 222The objects of interest in CTA are on average two orders of magnitude larger in relation to the size of the tomogram than those in cryoET, where the particles often span only tens of voxels., it was not clear whether a similar NST based framework could be successfully applied in cryoET to image nanoscale particles. In addition, the article does not provide source code and does not document the experiments in enough detail for us to be able to reproduce the results or adapt the method to the cryoET domain.
@@ 1.4 Other simulators
The value of simulated data in cryoEM is well recognized and micrograph simulation has been attempted several times. In first approximation a single cryoEM micrograph is the projection of a 3D object, convolved with the electron microscope’s point spread function. Additionally, the overall process comprises several sources of noise coming from the nature of the sample, the microscope, and the imaging procedure which are hard to model accurately. For a better overview of the attempts, it is necessary to mention TEM simulators that were developed in the past decades to simulate micrographs in cryo conditions.
@@
Earlier works provided fast simplistic models mostly based on additive white Gaussian noise333This motivated our choice of baseline. or coloured noise Scheible et al. (2021). Other works originated from the insights into the physics of TEM image formation and advanced the simulations by modeling various sources of noise, e.g. an improvement in modeling the structural noise was proposed in TEM Simulator (C) Rullgård et al. (2011). Another improvement was done in InSilicoTEM (Matlab) presented in Vulović et al. (2013) by implementing the multislice method originally proposed in Cowley and Moodie (1957). This makes it the next most relevant simulator related to our work after SHREC which is, as many others, also based on the same multislice method 444This motivated our choice of benchmark.. To the best of our knowledge, the most recent improvement presented in Himes and Grigorieff (2021) is distributed as a part of cisTEM package (C++). It introduced frozen plasmon method to explicitly model spatially variable inelastic scattering processes in cryo-electron microscopy. The aforementioned works, based on the same multislice method, however, suffer from a heavy computational burden or are limited to simulations of a single molecular complex. Despite these limitations, they represent a set of precise advanced physics-based simulators and as such also provide a comprehensive literature survey referring the reader to a body of detailed resources about modeling the image formation in transmission electron microscopy.
In material sciences, GPU accelerated simulators such as MULTEM (C++, CUDA) Lobato and Van Dyck (2015); Lobato et al. (2016), abTEM (Python) Madsen and Susi (2020), or Prismatic (C++, CUDA) Pryor et al. (2017) have emerged. However, as reviewed in Kirkland (2020), advanced TEM simulators used in material sciences require atomic models of the background and the entire specimen. Moreover, in cryoEM, the sample and its interaction with the electron beam is fundamentally different. While in material science samples are often only a few layers of atoms thick and the imaged atoms have strong interaction with the electron beam, cryoEM samples are thousands of atom layers thick, are less ordered, and only weakly interact with the electron beam. Additionally, cryoEM samples can only withstand little radiation before complete destruction resulting in much lower signal to noise ratios. Due to these differences, such simulators are of limited use for macromolecular biological specimen simulations that we experiment with in this article. Finally, in contrary to our approach, all aforementioned methods require precise calibration protocols for setting the values of simulation parameters.
@@ 1.5 Structure of this article
To evaluate the methods, we used the simulated data by each of them as a training set for the DeepFinder neural network (more in Section 4.1). The network was trained to solve two evaluation tasks proposed in SHREC challenge. The final comparison was done by observing the models’ performances on the test tomogram chosen by SHREC. Detailed description of the evaluation is in Section 4. More on experiments and results in Section 5. This paper is accompanied with source code, the results are fully-reproducible, and the full experiment results are available in the repository.
Side-by-side comparison of all projections used in this article is available in Appendix A. The process of re-creating the benchmark data is detailed in Appendix B. Fine-tuning experiments that further improved the DF performance are described in Appendix C. Additional experiments, NST ablation study, and investigation of DF limits are presented in Appendix D. Further details on the performance of studied models on the particle classification task are provided in Appendix E.
@@ 2 Data
For the purposes of this paper, we used the latest version of the SHREC data set from the year 2021. More specifically, we based the presented methods on full-resolution grandmodels (synthetic volumes containing randomly scattered particles) and simulated projections from the SHREC 2021 additional data set that was made publicly available555Download from shrec.net/cryo-et or doi:10.34894/XRTJMA.. An example visualization of a simulated projection using the SHREC simulator is depicted in Figure 1. We use the unchanged simulated projections as a benchmark which we try to approximate with our proposed method.
@@
The steps to create the simulations were described in Gubins et al. (2021). However, at the time of writing of this paper, the implementation is not publicly available and from the description of the steps it is clear that the method is, and rightfully so, very elaborate, and not at all elementary to reimplement (also due to various steps being under-documented, see Appendix B).
@@
On the following lines, we briefly summarize the steps needed to create the SHREC data set. The authors first constructed 3D ground-truth specimens (grandmodels) along with annotations. Each specimen contained uniformly distributed and rotated protein instances (represented as Coulomb density) of varying size and structure from the Protein Data Bank (PDB) Berman et al. (2000), as well as membranes and gold fiducials, which are commonly found in tomograms. Then, they simulated a layer of amorphous ice into the grandmodels before rotating them over 61 evenly-spaced tilt angles, ranging from -60∘ to +60∘, in order to be projected. The noiseless projections were produced using their own implementation of a TEM simulator based on the multislice approach presented in Vulović et al. (2013). Next, they sampled from a Poisson distribution with a specific electron dose to obtain the final electron counts (files labeled as projections). After that, they randomly shifted the projections to model the tilt misalignment and scaled the amplitudes in Fourier space using information about amplitudes from experimental images in order to increase their similarity (files labeled as projections_unbinned). In the end, they obtained the final images (files labeled as reconstruction) by reconstructing the binned projections using weighted-backprojection algorithm from a private version of the PyTom package Hrabe et al. (2012).
@@
We chose the SHREC data set as it is established, well-executed, downloadable, and allows researchers to compare their results with previous works without the need to reproduce all results from scratch. In this paper, we are not focusing on solving the proposed tasks, instead, we are using them as evaluation metrics to measure the performance of our proposed method. However, the availability of the data and descriptions of the methods made this research feasible.
@@
As stated earlier, we used the full-resolution grandmodels from the SHREC data set to create our own noiseless projections using Radon transform, as implemented in the scikit-image library for image processing in Python Van der Walt et al. (2014). We computed our own noiseless projections for three primary reasons. Firstly, SHREC only supplied full-resolution noiseless projections embedded within the simulated ice layer, a feature we also aimed to incorporate in our surrogate. Secondly, we intended to provide access to all implementation steps within the code-base. Lastly, it was necessary to produce noiseless reconstructions that could later be used to estimate the performance boundaries of DF on this data set. We also had to omit the usage of the reconstructions provided by SHREC because the exact configuration of all the steps is not published and the version of the PyTom software package used to create this data is not public. Also, the public version is not yet properly documented or straightforward to use. Thus, the only feasible option to ensure comparable results across our experiments was to create our own reconstructions from all the newly created projections described in Section 3. Nevertheless, we put a lot of effort into matching our reconstructions with the ones originally published in SHREC. We created them using the RadonTea Python package Müller, Paul (2013) using a custom filter described in Appendix B. For additional flexibility, we also implemented support for tomographic reconstructions using the well-known IMOD package Mastronarde and Held (2017). To ensure comparability also with previously published results, we used the original SHREC model_9 reconstruction as a test tomogram for all presented methods as it was done in all the past challenges.